Find three consecutive even integers such that five times the second equals twice the sum of the first and the third?
Answer by checkley77(12569)

LET X, X+2 & X+4 BE THE THREE EVEN INTEGERS.
5(X+2)=2(X+X+4)
5X+10=2(2X+4)
5X+10=4X+8
5X-4X=8-10
X=-2 IS THE FIRST INTEGER.
-2=2=0 IS THE SECOND INTEGER.
-2+4=2 IS THE THIRD INTEGER.
PROOF:
5*0=2(-2+2)
0=2*0
0=0